OpenStudy (anonymous):

simplify the expression 10x^5 y^3/ 2x^6y

OpenStudy (anonymous):

Just cancel the like terms.\[\frac{10x^5y^3}{2x^6y}=\frac{5y^2}{x}\]

OpenStudy (anonymous):

More correctly, cancel the factors in common.